OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (anonymous):

y = -4x+5

OpenStudy (lalaly):

since its perpindicular the slope is the negative reciprocal of 4 = \[\frac{-1}{4}\] so the equation y=mx+b \[-3=\frac{-1}{4}(2)+b\]\[-12=-2+b\]\[b=-10\] now we have \[y=\frac{-1}{4}x-10\] in standard form it is \[x+4y+40=0\]

OpenStudy (anonymous):

oooops, i forgot the reciprocal, my bad,

OpenStudy (nilankshi):

hmmmmmm